OpenStudy (anonymous):

find the derivative of f(x) = -3/(-4x+5)

OpenStudy (anonymous):

Have you tried the Quotient Rule?

OpenStudy (anonymous):

i wouldn't use it \[\left(\frac{1}{f}\right)'=-\frac{f'}{f^2}\]